# mO Xiusi[#1]

### v1

@@ 4.4 Conversational Recommendation
Building upon the previous sections where knowledge injection was performed using textual contexts, we explore injecting conversational data into the model. Specifically, we adopt the Conversational Recommendation task, which involves multi-turn interactions between users and a recommendation system. The intuition is that exposure to such conversations enables the model to generate more accurate and user-aligned movie recommendations. We utilize two datasets: (1) INSPIRED (Hayati et al., 2020): Contains 731 conversational interactions. (2) REDIAL (Li et al., 2018): Comprises 7,415 conversational interactions. Following the evaluation procedure from He et al. (2023), we prompt the target instruct model to generate 20 movie recommendations based on user queries. The recommendations are then evaluated using recall metrics ($r_{1},r_{2},r_{3},r_{4}$), detailed in Appendix A.1.1. Our task formulation involves injecting a substantial number of conversations (731 for INSPIRED and 1,000 for REDIAL) into the model’s parameters using SELF-PARAM, aiming to enhance recommendation quality compared to the backbone model without access to these conversations. We use the Mistral-7B-instruct-v0.2 model as our backbone. We compare SELF-PARAM against the following baselines: (1) FT (C): Fine-tuning the model on conversations, with loss calculated solely on system utterances; (2) FT (Q): Fine-tuning the model on generated question-answer pairs, with loss calculated only on answers; (3) Dense Passage Retrieval (DPR) (Karpukhin et al., 2020): Encodes conversations and queries into embeddings, retrieving relevant conversations using faiss; (4) BM25: A sparse retriever-based method for retrieving relevant conversations; (5) RAPTOR (Sarthi et al., 2024): Constructs knowledge graphs from conversations using gpt-4o-mini and retrieves relevant information from the graph. For retrieval-augmented generation (RAG) methods (DPR, BM25, RAPTOR), all conversations are stored in a knowledge pool. During inference, the most relevant conversations are retrieved and incorporated into the prompt as context.
@@
The results are presented in Table 3. SELF-PARAM consistently outperforms all baselines across both datasets, demonstrating its effectiveness in enhancing recommendation quality without requiring additional storage modules. From Table 3, the following observations can be made: (1) SELF-PARAM achieves the highest Recall@1 scores across all scenarios and datasets, outperforming all baselines. This underscores the effectiveness of injecting conversations directly into the model parameters, enabling better understanding and recommendation generation. (2) Both FT (C) and FT (Q) exhibit zero performance across all scenarios, indicating that traditional fine-tuning methods fail to effectively integrate conversational knowledge. This may be due to the divergent styles between recommendation conversations and the original instruct models, leading to degradation in model behavior. However, unlike fine-tuning baselines, SELF-PARAM maintains consistent performance without disrupting the model’s inherent capabilities. This is attributed to the use of KL divergence in our training objective, which aligns the target model’s behavior with the backbone model while integrating new knowledge. (3) Methods like DPR, BM25, and RAPTOR, which rely on external retrieval mechanisms, achieve little or no improvements. This may be due to the difficulty in effectively retrieving and integrating relevant conversations from a large pool, limiting their ability to enhance recommendation quality. Overall, SELF-PARAM demonstrates significant improvements in recommendation accuracy by effectively internalizing conversational knowledge without additional storage overhead. This highlights the potential of SELF-PARAM in enhancing interactive and dynamic tasks within large language models.
5 Conclusion
@@
In this paper, we addressed the critical challenge of integrating small-scale experiences into large language models (LLMs) with high efficacy and long-term retention without incurring additional storage complexity. We introduced SELF-PARAM (Self-Updatable Large Language Models with Parameter Integration), a novel method that embeds experiences directly into model parameters by minimizing the Kullback-Leibler (KL) divergence between an original model with context and a target model without context. Through comprehensive evaluations across diverse tasks, including Single Context Injection, Batch Context Injection, Sequential Injection, and Conversational Recommendation, we find that SELF-PARAM consistently outperformed existing baselines, demonstrating superior performance in both immediate knowledge integration and robust retention over multiple injections. Notably, SELF-PARAM achieved these advancements without requiring additional parameters or external storage modules, thereby maintaining the model’s integrity and scalability in dynamic environments. These findings highlight the potential of SELF-PARAM to enhance the adaptability and efficiency of LLMs. Future work may explore scaling SELF-PARAM to larger models, incorporating multi-modal experiences, and applying the method to a broader range of application domains to further validate its effectiveness and versatility.
@@ A.3 Construction of Target Sentence Set
The prompt used for query the instruct model using the context is shown below:

### v2

@@ 4.4 Conversational Recommendation
Building upon the previous sections where knowledge injection was performed using textual contexts, we explore injecting conversational data into the model. Specifically, we adopt the Conversational Recommendation task, which involves multi-turn interactions between users and a recommendation system. The intuition is that exposure to such conversations enables the model to generate more accurate and user-aligned movie recommendations. We utilize two datasets: (1) INSPIRED (Hayati et al., 2020): Contains 731 conversational interactions. (2) REDIAL (Li et al., 2018): Comprises 7,415 conversational interactions. Following the evaluation procedure from He et al. (2023), we prompt the target instruct model to generate 20 movie recommendations based on user queries. The recommendations are then evaluated using recall metrics ($r_{1},r_{2},r_{3},r_{4}$), detailed in Appendix A.1.1. Our task formulation involves injecting a substantial number of conversations (731 for INSPIRED and 1,000 for REDIAL) into the model’s parameters using SELF-PARAM, aiming to enhance recommendation quality compared to the backbone model without access to these conversations. We use the Mistral-7B-instruct-v0.2 model as our backbone. We compare SELF-PARAM against the following baselines: (1) FT (C): Fine-tuning the model on conversations, with loss calculated solely on system utterances; (2) FT (S): Fine-tuning the model on generated question-answer pairs, with loss calculated only on answers; (3) Dense Passage Retrieval (DPR) (Karpukhin et al., 2020): Encodes conversations and queries into embeddings, retrieving relevant conversations using faiss; (4) BM25: A sparse retriever-based method for retrieving relevant conversations; (5) RAPTOR (Sarthi et al., 2024): Constructs knowledge graphs from conversations using gpt-4o-mini and retrieves relevant information from the graph. For retrieval-augmented generation (RAG) methods (DPR, BM25, RAPTOR), all conversations are stored in a knowledge pool. During inference, the most relevant conversations are retrieved and incorporated into the prompt as context.
@@
The results are presented in Table 3. SELF-PARAM consistently outperforms all baselines across both datasets, demonstrating its effectiveness in enhancing recommendation quality without requiring additional storage modules. From Table 3, the following observations can be made: (1) SELF-PARAM achieves the highest Recall@1 scores across all scenarios and datasets, outperforming all baselines. This underscores the effectiveness of injecting conversations directly into the model parameters, enabling better understanding and recommendation generation. (2) Both FT (C) and FT (S) exhibit zero performance across all scenarios, indicating that traditional fine-tuning methods fail to effectively integrate conversational knowledge. This may be due to the divergent styles between recommendation conversations and the original instruct models, leading to degradation in model behavior. However, unlike fine-tuning baselines, SELF-PARAM maintains consistent performance without disrupting the model’s inherent capabilities. This is attributed to the use of KL divergence in our training objective, which aligns the target model’s behavior with the backbone model while integrating new knowledge. (3) Methods like DPR, BM25, and RAPTOR, which rely on external retrieval mechanisms, achieve little or no improvements. This may be due to the difficulty in effectively retrieving and integrating relevant conversations from a large pool, limiting their ability to enhance recommendation quality. Overall, SELF-PARAM demonstrates significant improvements in recommendation accuracy by effectively internalizing conversational knowledge without additional storage overhead. This highlights the potential of SELF-PARAM in enhancing interactive and dynamic tasks within large language models.
4.5 Ablation Study
Ablation Study of Training Objective
Since we introduce a new training objective, as described in Section 3.1.1, it is also important to examine the performance when applying Next-Word-Prediction loss directly to either the original context or the constructed target sentence set. To explore this, we include two baseline settings, FT (C) and FT (S), in Tables 1, 2, and 3, which correspond to fine-tuning on the original contexts and the target sentence set, respectively. The results show that while fine-tuning on the target sentence set (FT (S)) generally outperforms fine-tuning on the original contexts (FT (C)), it still falls significantly short of our proposed approach. This highlights the importance of our training objective in achieving superior performance.
Ablation Study of the Model for Target Sentence Set Construction
100 Contexts 500 Contexts
Mistral-7B Llama3-8B Mistral-7B Llama3-8B
4o-mini 0.4521 0.4368 0.4384 0.4221
instruct 0.4502 0.4341 0.4836 0.4464
Table 4: Ablation Study of the Model For Target Sentence Set Construction. Here 4o-mini refers to gpt-4o-mini, instruct means using the corresponding instruct model, i.e., using Mistral-7B-Instruct-v0.3 for Mistral-7B and Llama-3-8B-Instruct for Llama3-8B.
To demonstrate that our performance gains do not stem from using a stronger model – and to distinguish our approach from knowledge distillation methods that transfer knowledge from larger models to smaller ones – we conduct an experiment where we replace the Target Sentence Set generated by gpt-4o-mini (as described in Section 4.2) with a corresponding instruct model. This ensures that no additional knowledge from a more advanced model is introduced, aligning with our ”self-updatable” objective. The results, presented in Table 4, confirm that our method remains effective even without leveraging a more powerful model to construct the target sentence set. This further validates the robustness of our approach.
5 Conclusion and Future Work
@@
In this paper, we addressed the critical challenge of integrating small-scale experiences into large language models (LLMs) with high efficacy and long-term retention without incurring additional storage complexity. We introduced SELF-PARAM (Self-Updatable Large Language Models with Parameter Integration), a novel method that embeds experiences directly into model parameters by minimizing the Kullback-Leibler (KL) divergence between an original model with context and a target model without context. Through comprehensive evaluations across diverse tasks, including Single Context Injection, Batch Context Injection, Sequential Injection, and Conversational Recommendation, we find that SELF-PARAM consistently outperformed existing baselines, demonstrating superior performance in both immediate knowledge integration and robust retention over multiple injections. Notably, SELF-PARAM achieved these advancements without requiring additional parameters or external storage modules, thereby maintaining the model’s integrity and scalability in dynamic environments. These findings highlight the potential of SELF-PARAM to enhance the adaptability and efficiency of LLMs. Future work may explore scaling SELF-PARAM to larger models, incorporating more diverse types of experiences, and applying the method to a broader range of applications such as sessions of conversations.
Ethics Statement
In this research, we use publicly available datasets (PwC, INSPIRED, and REDIAL) that do not contain personally identifiable information, ensuring compliance with data privacy standards and licensing agreements. By embedding contextual knowledge directly into model parameters, there is a potential risk of data leakage; to address this, we adhere to robust security practices and recommend further safeguards for deployment. All methodologies and data handling procedures comply with relevant legal frameworks and ethical guidelines, and we have maintained research integrity through transparent reporting and appropriate citations. Overall, while our approach enhances the capabilities of large language models, we remain committed to promoting fairness, security, and responsible innovation in artificial intelligence.
Reproducibility Statement
We describe construction process of the target sentence set in § A.3.
Single Context Injection (§ 4.1): We describe the dataset setup in § 4.1 and describe the implementation details in § A.2.
Batch Context Injection (§ 4.2): Similarly, we describe the dataset setup in § 4.2 and describe the implementation details in § A.2.
Sequential Injection (§ 4.3): The dataset setup is described in § 4.3 and the implementation details are shown in § A.2.
Conversational Recommendation (§ 4.4): The dataset setup is described in § 4.4 and the implementation details are described in § A.2.
@@ A.3 Construction of Target Sentence Set
The prompt used for querying the instruct model using the context is shown below:
@@
Appendix B Additional Experiments
B.1 Sensitivity to Target Sentence Set Diversity
Our training objective, Eq.(3), uses 50 question-answer pairs as the target sentence set per context over one epoch. During our exploration of this project, we found that training on a smaller set of 10 question-answer pairs for five epochs led to lower performance. We report the results we obtained in Table 6. This analysis suggests that question-answer diversity is important to ensure comprehensive attention to context information. Limited diversity may lead to slightly reduced performance.
Model & # of Contexts 50 QA pairs (1 Epoch) 10 QA pairs (5 epochs)
OpenLlama-3B-v2 & 100 0.5082 0.4203
OpenLlama-3B-v2 & 500 0.5048 0.4203
Mistral-7B & 100 0.4521 0.3891
Mistral-7B & 500 0.4384 0.3813
Table 6: Sensitivity to Target Sentence Set Diversity. Here we consider batch context injection.
Figure 3: Average QA-F1 scores after sequentially injecting contexts into the model across 50 sequences without SlimPajama
Figure 4: Average QA-F1 scores after sequentially injecting contexts into the model across 50 sequences with fine-tuning on the Target Sentence Set.
Figure 5: Average QA-F1 scores after sequentially injecting contexts into the model across 50 sequences with fine-tuning on the context.
B.2 Ablation Study for Sequential Injection
To conduct the ablation study for sequential injection, we consider the following settings:
  • SELF-PARAM w/o SlimPajama: As mentioned in Section A.3, we adopt SlimPajama as the training set to retain the model’s general abilities. Thus we conduct the experiments with exactly the same settings while skipping the steps of training on SlimPajama. The results are shown in Figure 3. Looking at the dianogal results [0.43, 0.57, 0.44, 0.52, 0.5, 0.67, 0.45, 0.49, 0.55, 0.48, 0.46, 0.4, 0.48, 0.35, 0.36, 0.34, 0.42, 0.24, 0.28, 0.27], we can see that the model’s generalization ability has be severely affected after around 20 steps of injection, where the model can only achieve less than 0.30 QA-F1-score when we are using Eq.(2) to inject new knowledge. In contrast, as shown in Figure 2, SELF-PARAM can maintain the QA-F1-score as always around 0.5. This justifies that we need SlimPajama to maintain the model’s general functionality.
  • Finetuning on the Context: Following the settings FT (C) in Section 4.1 and 4.2, at each step, we fine-tune the base model, Mistral-7B, with LoRA applied, on a single context for 50 epochs. Similarly, SlimPajama is used as the regularization set during fine-tuning. After completing each step, we merge the LoRA weights into the original model before proceeding to fine-tune on the next context. To evaluate performance, we construct 25 sequences, each consisting of 20 consecutive contexts, and report the average performance across all 25 sequences (These settings are consistent with those described in Section 4.3). The results are shown in Figure 4. From the figure, we can see that FT (C) demonstrates the lowest efficacy. By omitting the first column and examining the diagonal values as highlighted, it is evident that FT (C) barely injects knowledge into the model.
  • Finetuning on the Target Sentence Set: Following the settings FT (S) in Section 4.1 and 4.2, we fine-tune the base model on the constructed question-answering pairs (i.e., the target sentence set) constructed from gpt-4o-mini with the same experimental settings as SELF-PARAM, with the results shown in Figure 5. From the figure, we can find that FT (S) has a retention ability comparable to ours. As shown in the first row of Figure 2 and 5, which shows the accuracy of querying the model with the question related $x_{1}$ after injecting $x_{1},\cdots,x_{n},n=1,\cdots,20$ into the model (here $x_{1},\cdots,x_{20}$ refer to 20 contexts that are sequentially injected), Our approach starts at a QA-F1 score of around 0.5 after injecting x1 and drops to approximately 0.33 after injecting x20 (Figure 2). In contrast, FT (S) begins at 0.36 and decreases to 0.31. Despite the similar retention ability, FT (S) has much lower efficacy, as shown by the diagonal values of the figures (meaning the QA-F1 score when answering the question related to $x_{n}$ after injecting $x_{1},\cdots,x_{n}$ into the model).